OpenStudy (anonymous):

θ = 135° is a central angle that cuts off an arc of length s = 3 km. In this case, find the radius of the circle. (Round the answer to two decimal places.)

OpenStudy (anonymous):

ohh i forgot this stuff

OpenStudy (anonymous):

Oh, alright.

Directrix (directrix):

135/360 = 3/ (2 pi r)

OpenStudy (anonymous):

yes that is correct

OpenStudy (anonymous):

ummm maybe not

Directrix (directrix):

Finish it - I was just passing by.

OpenStudy (anonymous):

k wait i just gotta see what u did

Directrix (directrix):

Circumference of circle can be thought of as 360 degrees which in a sense is equivalent to the perimeter of the circle, 2 pi r.Set up a proportion.

OpenStudy (anonymous):

3*(360/135)=r r=8

Directrix (directrix):

r = 4/pi r = 1.27 km

OpenStudy (anonymous):

hehe i wonder who is right